# Dirac quantum criticality in twisted double bilayer transition metal dichalcogenides

### v1

@@ Abstract
We investigate the phase diagram of moiré double bilayer transition metal dichalcogenides with ABBA stacking as a function of twist angle and applied pressure. At hole filling $\nu=2$ per moiré unit cell, the noninteracting system hosts a Dirac semimetal with graphene-like low-energy bands in the moiré Brillouin zone. At small twist angles, the Fermi velocity is reduced and interactions dominate the low-temperature behavior. A strong-coupling analysis identifies insulating ferromagnetic and antiferromagnetic ground-state candidates, characterized by spin-density modulations set by the moiré scale. Using a realistic continuum model with long-range Coulomb interactions, we perform self-consistent Hartree-Fock calculations to study the competition between these states. Varying the twist angle or pressure drives a transition from a Dirac semimetal to an antiferromagnetic insulator, which breaks SU(2) spin rotation and two-fold lattice rotation symmetries. This semimetal-to-insulator transition is continuous and belongs to the (2+1)D relativistic Gross-Neveu-Heisenberg universality class with $N=2$ four-component Dirac fermions. Finite heterostrain, relevant in realistic samples, induces a crossover from Gross-Neveu-Heisenberg universality at intermediate temperatures to conventional (2+1)D Heisenberg criticality at the lowest temperatures. Further decreasing the twist angle can cause a level crossing from the antiferromagnetic insulator into a ferromagnetic insulator with spin-split bands. Our results provide a comprehensive theoretical framework that complements and elucidates recent experiments in twisted double bilayer WSe2.
@@ I Introduction
Quantum critical points involving Dirac fermions arise across multiple domains, bridging phenomena between condensed-matter and high-energy physics. They have been studied in the context of graphene [2, 3, 4, 5, 6, 7, 8, 9], topological insulators [10, 11], frustrated quantum magnets [12, 13, 14, 15, 16, 17], and $d$-wave superconductors [18, 19, 20], as well as in toy models of chiral symmetry breaking and nonperturbative renormalizability [21, 22, 23, 24, 25, 26, 27, 28]. In such systems, gapless Dirac fermions can acquire a dynamical mass gap through interactions, giving rise to nontrivial quantum critical behavior that has been extensively explored theoretically [29, 30]. For a long time, this physics remained purely theoretical and beyond experimental reach. Very recently, however, it has become possible to realize and probe Dirac quantum critical points in moiré materials tuned by twist angle. Theoretical studies of twisted bilayer graphene predicted a continuous transition between a Dirac semimetal and a Kramers intervalley-coherent insulator [31, 32], belonging to the Gross-Neveu-XY universality class [33, 34]. Concurrently, experiments on twisted double bilayer WSe2 observed a Dirac-semimetal-to-insulator transition, interpreted as a realization of a related type of Dirac quantum critical point [35].
@@
In this work, we theoretically investigate the phase diagram of hole-doped double bilayer transition metal dichalcogenides (TMDs) with a small twist angle between the two bilayers, see Fig. 1(a). TMDs are layered materials with chemical formula $MX_{2}$, where $M$ is a transition metal, taken here to be a group-VI element such as Mo or W, and $X$ is a chalcogen such as S, Se, or Te. Earlier studies have placed particular emphasis on bilayer TMDs [36, 37, 38, 39, 40, 41, 42]. In such systems, however, the location of the valence-band maximum depends sensitively on the specific choice of metal and chalcogen [43, 44, 45], making it challenging to realize a Dirac quantum critical point with SU(2) spin symmetry due to the spin splitting induced by spin-orbit coupling. In contrast, the stronger interlayer hybridization in trilayer and double bilayer TMDs is generally expected to pin the valence-band maximum at the Brillouin-zone center $\boldsymbol{\Gamma}$ [46, 47], analogous to the situation in bulk TMDs [48]. The bands near $\boldsymbol{\Gamma}$ are spin degenerate, giving rise to an SU(2)-symmetric effective low-energy description. Here we focus on ABBA-stacked double bilayer TMDs. Introducing a small twist between the two bilayers generates a tunable emergent honeycomb lattice that retains the SU(2) spin symmetry and two-fold lattice rotation symmetry of monolayer graphene, now realized on the moiré scale [49]. Note that the crystal structure of ABBA-stacked double bilayer TMDs [49, 35] differs from that of ABAB-stacked double bilayers [47], even when a finite twist is applied between the two inner layers, due to the absence of the two-fold rotation symmetry $C_{2y}$ about the in-plane $y$ axis indicated in Fig. 1(a). In the noninteracting limit, the electronic spectrum hosts two gapless Dirac cones located at the moiré Brillouin-zone corners $\boldsymbol{\kappa}$ and $\boldsymbol{\kappa}$’, see Fig. 1(b). At a hole filling of $\nu=2$ per moiré unit cell, the Fermi level sits precisely at the band crossing, with no additional low-energy electronic states away from the $\boldsymbol{\kappa}$ and $\boldsymbol{\kappa}$’ points. To investigate the low-temperature behavior of the system, we use a realistic continuum model with long-range Coulomb interactions, employing parameters relevant to experiments on double bilayer WSe2 [35]. Our model consistently incorporates the angle dependence of the interaction strength and includes contributions from remote bands. It also enables the study of uniaxial pressure and residual heterostrain, as expected in realistic samples. At large twist angles, the Dirac semimetal remains stable for realistic interactions. Decreasing the twist angle, or applying uniaxial pressure at sufficiently small angles, destabilizes the semimetal and induces a transition to an antiferromagnetic insulator with broken SU(2) spin and inversion symmetries, featuring a spin-density wavelength on the scale of the moiré lattice constant. The semimetal-to-insulator transition is continuous and falls within the (2+1)D relativistic Gross-Neveu-Heisenberg universality class with $N=2$ four-component Dirac fermions. Finite heterostrain drives a crossover from Gross-Neveu-Heisenberg universality at intermediate temperatures to conventional (2+1)D Heisenberg criticality at ultralow temperatures. At very small twist angles, the antiferromagnet competes with an insulating state exhibiting a finite total magnetization while preserving inversion symmetry. For sufficiently strong Coulomb interactions, decreasing the twist angle can drive a second transition from the antiferromagnetic insulator to this ferromagnetic insulator. This antiferromagnet-to-ferromagnet transition is found to be discontinuous. While our calculations are performed using a model relevant for double bilayer WSe2, we expect the qualitative features of our results to apply more broadly to twisted TMDs that host an emergent honeycomb lattice on the moiré scale.
@@ II.5 Residual heterostrain
One of the most significant perturbations in moiré heterostructures is residual heterostrain, which corresponds to a relative in-plane distortion between adjacent layers. Even small amounts of heterostrain can substantially modify the moiré band structure and symmetry properties, and since it is generally expected to occur in realistic samples, it must be incorporated in any realistic analysis [66, 67, 68]. To illustrate the impact of strain on twisted double bilayer TMDs, we focus on uniaxial heterostrain applied between the two bilayers. Because the low-energy physics is dominated by the $\boldsymbol{\Gamma}$ valley, the quantitative effect of heterostrain on the electronic band structure is relatively small. Nonetheless, it breaks the $C_{2y}$ symmetry of the double bilayer, leading to notable consequences for the critical behavior at ultralow temperatures.
In $\mathbf{K}$-valley moiré heterostructures, strain generally has a twofold effect: it modifies the lattice geometry and shifts the microscopic $\mathbf{K}$ and $\mathbf{K}^{\prime}$ valleys due to anisotropic changes in the in-plane bond lengths, typically captured via minimal coupling to a phenomenological vector potential [69, 70, 67]. In contrast, an analogous shift of the $\boldsymbol{\Gamma}$-valley in the double bilayer TMDs studied here is forbidden, as it would violate the $C_{2y}$ symmetry of the individual layer. Consequently, we only need to account for the strain-induced geometric changes. Importantly, while heterostrain preserves the $C_{2y}$ symmetry of individual layers, it breaks both the $C_{2y}$ and $C_{3z}$ symmetries of the ABBA-stacked double bilayer.
@@ IV Strong-coupling limit
By analogy with the strong-coupling limit of twisted bilayer graphene, we begin by assuming an approximate chiral symmetry $\mathcal{C}$ in the noninteracting spectrum. This assumption is justified at small twist angles near $1^{\circ}$, where the bandwidth of the two spin-degenerate low-energy bands is narrow [cf. Fig. 3(f)], even though no parameter choice of the continuum model renders the symmetry exact. The chiral limit nevertheless provides a useful framework for building intuition about the competing orders at small twist angles. Antichiral terms present in the realistic model are then included as perturbative corrections.
@@
where $\mathcal{K}$ denotes complex conjugation, $\sigma_{x}$ and $\sigma_{z}$ are Pauli matrices acting on the sublattice degrees of freedom, and the Pauli matrix $s_{y}$ acts on the spin degree of freedom. Note that $C_{2z}$ and $\mathcal{T}$ invert wavevectors, whereas $\mathcal{C}$ leaves them unchanged. Decomposing the overlap matrix into symmetric and antisymmetric parts under chiral transformations, $\Lambda(\mathbf{k},\mathbf{q})=\Lambda_{\mathrm{S}}(\mathbf{k},\mathbf{q})+\Lambda_{\mathrm{A}}(\mathbf{k},\mathbf{q})$, with $\Lambda_{\mathrm{S}/\mathrm{A}}(\mathbf{k},\mathbf{q})=\frac{1}{2}\left(\Lambda(\mathbf{k},\mathbf{q})\pm\sigma_{z}\Lambda(\mathbf{k},\mathbf{q})\sigma_{z}\right)$ enforces the forms
@@
Figure 2: (a) Hartree-Fock energy $E_{\mathrm{HF}}$ of the antiferromagnetic (blue) and ferromagnetic (red) states relative to the noninteracting ground-state energy $E_{0}$ in the full model [Eq. (5)] for an effective permittivity $\epsilon_{\mathrm{eff}}=110$, representative of realistic values for twisted double bilayer WSe2. Calculations were performed on an $18\times 18$ momentum-space grid, keeping two bands per spin species, with initial states preserving the respective symmetries of each candidate phase. At small twist angles, the two states are nearly degenerate, reflecting the approximate chiral symmetry, yet the antiferromagnetic state is ultimately favored across the entire range. (b) Same as (a) but for $\epsilon_{\mathrm{eff}}=10$, enhancing the relative weight of the antichiral interaction compared with the kinetic contribution. In this case, the ferromagnetic state is favored throughout the range of twist angles considered.
@@
for the antiferromagnetic state, with the latter being strictly positive if there are antichiral contributions to the overlap matrices. Therefore, when the antichiral interaction is the dominant perturbation, the ferromagnetic state is energetically preferred.
To illustrate the competition between the two magnetic orders in the presence of both antichiral and kinetic contributions, Fig. 2 shows numerical results from a Hartree-Fock analysis of the full interacting continuum model as a function of twist angle $\theta$ for two different effective permittivities, $\epsilon_{\text{eff}}=110$ and $\epsilon_{\text{eff}}=10$. The relative energy difference between the two magnetic states decreases with decreasing twist angle, reflecting the approximate chiral symmetry at small angles. For the larger permittivity $\epsilon_{\text{eff}}=110$, which is of the order of realistic values, the antiferromagnetic state is favored because the kinetic contribution dominates over the antichiral terms, as shown in Fig. 2(a). By contrast, reducing the effective permittivity enhances the relative strength of the antichiral interaction, stabilizing the ferromagnetic state over a range of twist angles, as illustrated for $\epsilon_{\text{eff}}=10$ in Fig. 2(b). This interplay between antichiral interactions and the kinetic contribution underlies the appearance of both ferromagnetic and antiferromagnetic regions in the phase diagram discussed in the following section.
@@ V Angle-tuned Phase diagram
Figure 3: (a) Charge density $n(\mathbf{R})$ (grayscale) and spin density $\langle\hat{s}_{z}\rangle(\mathbf{R})$ (overlayed color scale) in the Dirac semimetal (DSM) ground state of the full Hamiltonian [Eq. (5)] at twist angle $\theta=4^{\circ}$ and effective permittivity $\epsilon_{\text{eff}}=110$, representative of realistic values for twisted double bilayer WSe2. Light gray lines indicate the microscopic lattices of the inner TMD layers. The charge density forms an emergent honeycomb pattern on the moiré scale. Densities are normalized to the value $n_{\mathrm{A}}$ at the A site of the emergent honeycomb lattice. (b) Same as (a), but for twist angle $\theta=2^{\circ}$. The ground state is antiferromagnetic (AFM), with antiparallel spins on the two sites of the emergent honeycomb lattice. The moiré scale is enlarged compared to (a). (c) Same as (a), but for twist angle $\theta=1^{\circ}$ and reduced effective permittivity $\epsilon_{\text{eff}}=40$. The ground state is ferromagnetic (FM), with parallel spins on the two sites of the emergent honeycomb lattice. (d) Electronic spectrum at twist angle $\theta=4^{\circ}$ and effective permittivity $\epsilon_{\text{eff}}=110$, representative of realistic values for twisted double bilayer WSe2. Light gray lines indicate the microscopic lattices of the inner TMD layers. Dirac points at the moiré Brillouin zone corners $\boldsymbol{\kappa}$ and $\boldsymbol{\kappa}$’ remain gapless and spin degenerate. Dashed gray lines show the noninteracting bands for comparison. Energies are referenced to the Fermi level $\varepsilon_{\mathrm{F}}$ at filling $\nu=2$ holes per moiré unit cell. (e) Same as (d), but for twist angle $\theta=2^{\circ}$. In contrast to the Dirac semimetal, the spectrum is fully gapped while remaining spin-degenerate, realizing an antiferromagnetic insulator. (f) Same as (d), but for twist angle $\theta=1^{\circ}$ and reduced effective permittivity $\epsilon_{\text{eff}}=40$. The spectrum is fully gapped and spin-split, realizing a ferromagnetic insulator. The inset indicates the path through the moiré Brillouin zone along which the energy bands are shown.
@@
We reiterate that for comparison with experiments at small twist angles, values of $\epsilon_{\text{eff}}\sim\mathcal{O}(100)$ are realistic [57, 31, 61]. Since $\epsilon_{\text{eff}}$ is dominated by internal screening, its precise value depends on $\theta$. Although a full calculation lies beyond the present work, we estimate $\epsilon_{\text{eff}}$ for twisted double bilayer WSe2 by matching the extrapolated Hartree-Fock gap $\Delta$ to experimental data from Ref. [35], yielding $\epsilon_{\text{eff}}(1.9^{\circ})\simeq 51$, $\epsilon_{\text{eff}}(2.5^{\circ})\simeq 109$, and $\epsilon_{\text{eff}}(2.7^{\circ})\simeq 112$, the latter corresponding to the experimentally observed critical angle $\theta_{\mathrm{c}}$ where the gap closes. These estimates suggest that the effective permittivity $\epsilon_{\text{eff}}$ increases with twist angle $\theta$ up to the critical angle $\theta_{\mathrm{c}}$, consistent with enhanced internal screening as the gap closes. For $\theta>\theta_{\mathrm{c}}$, $\epsilon_{\text{eff}}$ is expected to decrease with $\theta$, reflecting the widening of the low-energy bands. Figure 4(c) presents the quantum phase diagram as a function of twist angle $\theta$ and effective permittivity $\epsilon_{\text{eff}}$. At large twist angles and realistic $\epsilon_{\text{eff}}$, the Dirac semimetal is stable, while decreasing $\theta$ drives a continuous transition into the antiferromagnetic insulator. The semimetal-to-insulator phase boundary in Fig. 4(c) is obtained from a crossing-point analysis of the renormalization group invariant $R_{m}$, consistent with the order-parameter extrapolation in Fig. 4(a). At very small angles ($\sim 1^{\circ}$) and reduced $\epsilon_{\text{eff}}$, a ferromagnetic insulator emerges. The discontinuous antiferromagnet-to-ferromagnet boundary is determined from linear extrapolation of the energy offset $\Delta E=E_{\mathrm{HF}}[P]-E_{\mathrm{HF}}[P_{0}]$ across the transition, consistent with the level crossing in Fig. 4(b).

### v2

@@ Abstract
We investigate the phase diagram of moiré double bilayer transition metal dichalcogenides with ABBA stacking as a function of twist angle and applied pressure. At hole filling $\nu=2$ per moiré unit cell, the noninteracting system hosts a Dirac semimetal with graphene-like low-energy bands in the moiré Brillouin zone. At small twist angles, the Fermi velocity is reduced and interactions dominate the low-temperature behavior. A strong-coupling analysis identifies insulating ferromagnetic and antiferromagnetic ground-state candidates, characterized by spin-density modulations set by the moiré scale. Using a realistic continuum model with long-range Coulomb interactions, we perform self-consistent Hartree-Fock calculations to study the competition between these states. Varying the twist angle or pressure drives a transition from a Dirac semimetal to an antiferromagnetic insulator, which breaks SU(2) spin rotation and two-fold lattice rotation symmetries. Within a renormalization group analysis of the most general symmetry-allowed low-energy field theory, we show that this semimetal-to-insulator transition is continuous and belongs to the (2+1)D relativistic Gross-Neveu-Heisenberg universality class with $N=2$ four-component Dirac fermions. Finite heterostrain, relevant in realistic samples, induces a crossover from Gross-Neveu-Heisenberg universality at intermediate temperatures to conventional (2+1)D Heisenberg criticality at the lowest temperatures. Further decreasing the twist angle can cause a level crossing from the antiferromagnetic insulator into a ferromagnetic insulator with spin-split bands. Our results provide a comprehensive theoretical framework that complements and elucidates recent experiments in twisted double bilayer WSe2.
@@ I Introduction
Quantum critical points involving Dirac fermions arise across multiple domains, bridging phenomena between condensed-matter and high-energy physics. They have been studied in the context of graphene [2, 3, 4, 5, 6, 7, 8, 9], topological insulators [10, 11], frustrated quantum magnets [12, 13, 14, 15, 16, 17], and $d$-wave superconductors [18, 19, 20], as well as in toy models of chiral symmetry breaking and nonperturbative renormalizability [21, 22, 23, 24, 25, 26, 27, 28]. In such systems, gapless Dirac fermions can acquire a dynamical mass gap through interactions, giving rise to nontrivial quantum critical behavior that has been extensively explored theoretically [29, 30]. For a long time, this physics remained purely theoretical and beyond experimental reach. Very recently, however, it has become possible to realize and probe Dirac quantum critical points in moiré materials tuned by twist angle. Theoretical studies of twisted bilayer graphene predicted a continuous transition between a Dirac semimetal and a Kramers intervalley-coherent insulator [31, 32], belonging to the Gross-Neveu-XY universality class [33, 34]. At the same time, experiments on twisted double bilayer WSe2 [35] and twisted bilayer MoSe2 [36] have observed a Dirac-semimetal-to-insulator transition, interpreted as a realization of a related type of Dirac quantum critical point.
@@
In this work, we theoretically investigate the phase diagram of hole-doped double bilayer transition metal dichalcogenides (TMDs) with a small twist angle between the two bilayers, see Fig. 1(a). TMDs are layered materials with chemical formula $MX_{2}$, where $M$ is a transition metal, taken here to be a group-VI element such as Mo or W, and $X$ is a chalcogen such as S, Se, or Te. Earlier studies have placed particular emphasis on bilayer TMDs [37, 38, 39, 40, 41, 42, 43, 44]. In such systems, however, the location of the valence-band maximum depends sensitively on the specific choice of metal and chalcogen [45, 46, 47], making it challenging to realize a Dirac quantum critical point with SU(2) spin symmetry due to the spin splitting induced by spin-orbit coupling. In contrast, the stronger interlayer hybridization in trilayer and double bilayer TMDs is generally expected to pin the valence-band maximum at the Brillouin-zone center $\boldsymbol{\Gamma}$ [48, 49], analogous to the situation in bulk TMDs [50]. The bands near $\boldsymbol{\Gamma}$ are spin degenerate, giving rise to an SU(2)-symmetric effective low-energy description. Here we focus on ABBA-stacked double bilayer TMDs. Introducing a small twist between the two bilayers generates a tunable emergent honeycomb lattice that retains the SU(2) spin symmetry and two-fold lattice rotation symmetry of monolayer graphene, now realized on the moiré scale [51]. Note that the crystal structure of ABBA-stacked double bilayer TMDs [51, 35] differs from that of ABAB-stacked double bilayers [49], even when a finite twist is applied between the two inner layers, due to the absence of the two-fold rotation symmetry $C_{2y}$ about the in-plane $y$ axis indicated in Fig. 1(a). In the noninteracting limit, the electronic spectrum hosts two gapless Dirac cones located at the moiré Brillouin-zone corners $\boldsymbol{\kappa}$ and $\boldsymbol{\kappa}$’, see Fig. 1(b). At a hole filling of $\nu=2$ per moiré unit cell, the Fermi level sits precisely at the band crossing, with no additional low-energy electronic states away from the $\boldsymbol{\kappa}$ and $\boldsymbol{\kappa}$’ points. To investigate the low-temperature behavior of the system, we use a realistic continuum model with long-range Coulomb interactions, employing parameters relevant to experiments on double bilayer WSe2 [35]. Our model consistently incorporates the angle dependence of the interaction strength and includes contributions from remote bands. It also enables the study of uniaxial pressure and residual heterostrain, as expected in realistic samples. At large twist angles, the Dirac semimetal remains stable for realistic interactions. Decreasing the twist angle, or applying uniaxial pressure at sufficiently small angles, destabilizes the semimetal and induces a transition to an antiferromagnetic insulator with broken SU(2) spin and inversion symmetries, featuring a spin-density wavelength on the scale of the moiré lattice constant. The semimetal-to-insulator transition is continuous and falls within the (2+1)D relativistic Gross-Neveu-Heisenberg universality class with $N=2$ four-component Dirac fermions. Finite heterostrain drives a crossover from Gross-Neveu-Heisenberg universality at intermediate temperatures to conventional (2+1)D Heisenberg criticality at ultralow temperatures. At very small twist angles, the antiferromagnet competes with an insulating state exhibiting a finite total magnetization while preserving inversion symmetry. For sufficiently strong Coulomb interactions, decreasing the twist angle can drive a second transition from the antiferromagnetic insulator to this ferromagnetic insulator. This antiferromagnet-to-ferromagnet transition is found to be discontinuous. While our calculations are performed using a model relevant for double bilayer WSe2, we expect the qualitative features of our results to apply more broadly to twisted TMDs that host an emergent honeycomb lattice on the moiré scale.
@@ II.5 Residual heterostrain
One of the most significant perturbations in moiré heterostructures is residual heterostrain, which corresponds to a relative in-plane distortion between adjacent layers. Even small amounts of heterostrain can substantially modify the moiré band structure and symmetry properties, and since it is generally expected to occur in realistic samples, it must be incorporated in any realistic analysis [68, 69, 70]. To illustrate the impact of strain on twisted double bilayer TMDs, we focus on uniaxial heterostrain applied between the two bilayers. Because the low-energy physics is dominated by the $\boldsymbol{\Gamma}$ valley, the quantitative effect of heterostrain on the electronic band structure is relatively small. Nonetheless, it explicitly breaks the symmetries that protect the linear band crossings at the $\boldsymbol{\kappa}$ points of the moiré Brillouin zone, with important consequences for the critical behavior at ultralow temperatures.
In $\mathbf{K}$-valley moiré heterostructures, strain generally has a twofold effect: it modifies the lattice geometry and shifts the microscopic $\mathbf{K}$ and $\mathbf{K}^{\prime}$ valleys due to anisotropic changes in the in-plane bond lengths, typically captured via minimal coupling to a phenomenological vector potential [71, 72, 69]. In contrast, an analogous shift of the $\boldsymbol{\Gamma}$ valley in the double bilayer TMDs studied here vanishes within the present continuum model, as it would violate the $C_{2y}$ symmetry of each individual layer, which forms the starting point of our analysis. Consequently, we focus exclusively on the effects of strain-induced geometric changes. Importantly, while heterostrain preserves the $C_{2y}$ symmetry of the individual layers, it breaks this symmetry in the full continuum Hamiltonian by distorting the top and bottom bilayers differently. The threefold rotation symmetry $C_{3z}$ about the out-of-plane $z$ axis is already broken at the single-layer level. Effects beyond the present continuum model therefore do not introduce any additional symmetry breaking and are not expected to lead to further qualitative changes.
@@ IV Strong-coupling limit
By analogy with the strong-coupling limit of twisted bilayer graphene, we begin by assuming an approximate chiral symmetry $\mathcal{C}$ in the noninteracting spectrum. We demonstrate in Appendix B.3 that this assumption is justified at small twist angles near $1^{\circ}$, where the bandwidth of the two spin-degenerate low-energy bands is narrow [cf. Fig. 3(f)], even though no parameter choice of the continuum model renders the symmetry exact. The chiral limit nevertheless provides a useful framework for building intuition about the competing orders at small twist angles. Antichiral terms present in the realistic model are then included as perturbative corrections.
@@
where $\mathcal{K}$ denotes complex conjugation, $\sigma_{x}$ and $\sigma_{z}$ are Pauli matrices acting on the sublattice degrees of freedom, and the Pauli matrix $s_{y}$ acts on the spin degree of freedom. Note that $C_{2z}$ and $\mathcal{T}$ invert wavevectors, whereas $\mathcal{C}$ leaves them unchanged. Decomposing the overlap matrices into symmetric and antisymmetric parts under chiral transformations, $\Lambda(\mathbf{k},\mathbf{q})=\Lambda_{\mathrm{S}}(\mathbf{k},\mathbf{q})+\Lambda_{\mathrm{A}}(\mathbf{k},\mathbf{q})$, with $\Lambda_{\mathrm{S}/\mathrm{A}}(\mathbf{k},\mathbf{q})=\frac{1}{2}\left(\Lambda(\mathbf{k},\mathbf{q})\pm\sigma_{z}\Lambda(\mathbf{k},\mathbf{q})\sigma_{z}\right)$ enforces the forms
@@
Figure 2: (a) Hartree-Fock energy $E_{\mathrm{HF}}$ of the antiferromagnetic (blue) and ferromagnetic (red) states relative to the noninteracting ground-state energy $E_{0}$ in the full model [Eq. (5)] for an effective permittivity $\epsilon_{\mathrm{eff}}=110$, representative of realistic values for twisted double bilayer WSe2. Calculations were performed on an $18\times 18$ momentum-space grid, keeping two bands per spin species, with initial states preserving the respective symmetries of each candidate phase. At small twist angles, the two states are nearly degenerate, reflecting the approximate chiral symmetry, yet the antiferromagnetic state is ultimately favored across the entire range. (b) Same as (a) but for $\epsilon_{\mathrm{eff}}=40$, enhancing the relative weight of the antichiral interaction compared with the kinetic contribution. In this case, the ferromagnetic state is favored at sufficiently small twist angles.
@@
for the antiferromagnetic state, with the latter being strictly positive if there are antichiral contributions to the overlap matrices. Therefore, when the antichiral interaction is the dominant perturbation, the ferromagnetic state is energetically preferred. Similarly, we show in Appendix B.6 that the dispersion favors the antiferromagnetic state.
To illustrate the competition between the two magnetic orders in the presence of both antichiral and kinetic contributions, Fig. 2 shows numerical results from a Hartree-Fock analysis of the full interacting continuum model as a function of twist angle $\theta$ for two different effective permittivities, $\epsilon_{\text{eff}}=110$ and $\epsilon_{\text{eff}}=40$. The relative energy difference between the two magnetic states decreases with decreasing twist angle, reflecting the approximate chiral symmetry at small angles. For the larger permittivity $\epsilon_{\text{eff}}=110$, which is of the order of realistic values, the antiferromagnetic state is favored because the kinetic contribution dominates over the antichiral terms, as shown in Fig. 2(a). By contrast, reducing the effective permittivity enhances the relative strength of the antichiral interaction, stabilizing the ferromagnetic state over a range of twist angles, as illustrated for $\epsilon_{\text{eff}}=40$ in Fig. 2(b). This interplay between antichiral interactions and the kinetic contribution underlies the appearance of both ferromagnetic and antiferromagnetic regions in the phase diagram discussed in the following section.
@@ V Angle-tuned Phase diagram
Figure 3: (a) Charge density $n(\mathbf{R})$ (grayscale) and spin density $\langle\hat{s}_{z}\rangle(\mathbf{R})$ (overlaid color scale, opacity indicating magnitude) in the Dirac semimetal (DSM) ground state of the full Hamiltonian [Eq. (5)] at twist angle $\theta=4^{\circ}$ and effective permittivity $\epsilon_{\text{eff}}=110$, representative of realistic values for twisted double bilayer WSe2. Light gray lines indicate the microscopic lattices of the inner TMD layers. The charge density forms an emergent honeycomb pattern on the moiré scale. Densities are normalized to the value $n_{\mathrm{A}}$ at the A site of the emergent honeycomb lattice. In the Dirac semimetal phase, the spin density vanishes, rendering the overlaid color scale transparent. (b) Same as (a), but for twist angle $\theta=2^{\circ}$. The ground state is antiferromagnetic (AFM), with antiparallel spins on the two sites of the emergent honeycomb lattice. The moiré scale is enlarged compared to (a). (c) Same as (a), but for twist angle $\theta=1^{\circ}$ and reduced effective permittivity $\epsilon_{\text{eff}}=40$. The ground state is ferromagnetic (FM), with parallel spins on the two sites of the emergent honeycomb lattice. (d) Electronic spectrum at twist angle $\theta=4^{\circ}$ and effective permittivity $\epsilon_{\text{eff}}=110$, representative of realistic values for twisted double bilayer WSe2. Dirac points at the moiré Brillouin zone corners $\boldsymbol{\kappa}$ and $\boldsymbol{\kappa}$’ remain gapless and spin degenerate. Dashed gray lines show the noninteracting bands for comparison. Energies are referenced to the Fermi level $\varepsilon_{\mathrm{F}}$ at filling $\nu=2$ holes per moiré unit cell. (e) Same as (d), but for twist angle $\theta=2^{\circ}$. In contrast to the Dirac semimetal, the spectrum is fully gapped while remaining spin-degenerate, realizing an antiferromagnetic insulator. (f) Same as (d), but for twist angle $\theta=1^{\circ}$ and reduced effective permittivity $\epsilon_{\text{eff}}=40$. The spectrum is fully gapped and spin-split, realizing a ferromagnetic insulator. The inset indicates the path through the moiré Brillouin zone along which the energy bands are shown.
@@
Figure 4(c) presents the quantum phase diagram as a function of twist angle $\theta$ and effective permittivity $\epsilon_{\text{eff}}$. At large twist angles and realistic $\epsilon_{\text{eff}}$, the Dirac semimetal is stable, while decreasing $\theta$ drives a continuous transition into the antiferromagnetic insulator. The semimetal-to-insulator phase boundary in Fig. 4(c) is obtained from a crossing-point analysis of the renormalization group invariant $R_{m}$, consistent with the order-parameter extrapolation in Fig. 4(a). At very small angles ($\sim 1^{\circ}$) and reduced $\epsilon_{\text{eff}}$, a ferromagnetic insulator emerges. The discontinuous antiferromagnet-to-ferromagnet boundary is determined from linear extrapolation of the energy offset $\Delta E=E_{\mathrm{HF}}[P]-E_{\mathrm{HF}}[P_{0}]$ across the transition, consistent with the level crossing in Fig. 4(b).
We reiterate that for comparison with experiments at small twist angles, values of $\epsilon_{\text{eff}}\sim\mathcal{O}(100)$ are realistic [59, 31, 63]. In Appendix C, we estimate $\epsilon_{\text{eff}}$ in the Dirac semimetal phase using the random phase approximation [77]. For the range of twist angles $\theta$ considered here, $\epsilon_{\text{eff}}$ is dominated by internal screening and therefore exhibits a strong dependence on $\theta$. In the Dirac semimetal phase, $\epsilon_{\text{eff}}(\theta)$ increases upon reducing $\theta$, as the narrowing of the low-energy bands enhances electronic screening. For twisted double bilayer WSe2 at $\theta\simeq 2.7^{\circ}$, corresponding to the experimentally observed critical angle $\theta_{\mathrm{c}}$, we obtain $\epsilon_{\text{eff}}(2.7^{\circ})\simeq 82$, assuming a substrate permittivity of $\epsilon_{\text{env}}=5$. This estimate is in reasonable agreement with the semimetal-to-insulator phase boundary shown in Fig. 4(c), which occurs at a critical permittivity $\epsilon_{\text{eff},\mathrm{c}}\simeq 112$ for $\theta=2.7^{\circ}$. In the gapped phase, $\epsilon_{\text{eff}}(\theta)$ is expected to decrease upon reducing $\theta$, reflecting the opening of an excitation gap. Its magnitude can be inferred by matching the extrapolated Hartree-Fock gap $\Delta$ to the experimental data from Ref. [35], which yields $\epsilon_{\text{eff}}(2.5^{\circ})\simeq 109$ and $\epsilon_{\text{eff}}(1.9^{\circ})\simeq 51$, again in reasonable agreement with the expectation based on the random-phase-approximation estimates in the gapless phase. This suggests that the ferromagnetic phase may also be realized in twisted double bilayer WSe2 at small twist angles $\theta\lesssim 1^{\circ}$.